### v1

# Selection rule of canonical differential equations from Intersection theory

Abstract
Canonical differential equations matrix is the 1-$\mathrm{d}\log$-form coefficients of projecting ($n$+1)-$\mathrm{d}\log$-forms onto $n$-$\mathrm{d}\log$-forms master integrand. Intersection number with dual form in relative cohomology could achieve the projecting and give selection rules of canonical differential equations from the pole structure of $\mathrm{d}\log$ master integrand.
1 Introduction
Integration-By-Parts (IBP) [1] and IBP-based differential equation [2, 3, 4, 5] method are two of the most powerful methods in perturbation quantum field theory. When we appropriately select the master integrals, differential equations of them could be transformed into canonical form [6],
$\mathrm{d}f_{I}=\big{(}\mathrm{d}\Omega\big{)}_{IK}f_{K}\,,\quad\big{(}\mathrm{d}\Omega\big{)}_{IK}=\epsilon\sum_{i}\text{C}^{(i)}_{IK}~{}\mathrm{d}\log\text{W}^{(i)}(\bm{s})\,.$ (1)
in which $\big{(}\mathrm{d}\Omega\big{)}_{IK}$ is $\mathrm{d}\log$-form proportional to $\epsilon$. Thereby it enables analytical solving master integrals as multi-polylogarithm [7, 8]. This method is called Canonical Differential Equations (CDE) and has become the most important analytic method in Feynman integrals’ evaluation. The $\mathrm{d}\log\text{W}^{(i)}(\bm{s})$ in (1) are referred to as symbol letters, whose complete set is known as symbol alphabet, constrains the function space of analytic solutions and can be utilized for bootstrap such as [9, 10, 11, 12, 13]. Thus, the properties of symbol letters or says also the $\mathrm{d}\log\text{W}^{(i)}(\bm{s})$ in (1) are important and have been studied in [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
One way to "appropriately select the master integrals" to get CDE is constructing $\mathrm{d}\log$ integrand as master integrands (which is also related to leading singularity) [6, 33, 34, 35, 36, 37, 38]. In [35, 33, 34], people have shown Baikov representation [39] make the construction more convenient. Recently, [40] shows how CDE emerges from $\mathrm{d}\log$ integrand of master integrals via intersection theory [41, 42]. In that work, only leading-order (LO) contribution and next-to-leading-order (NLO) contribution of intersection number are needed and both have universal formulas and selection rules of CDE have been given. However, in [40], the analyzing needs to keep some regulators for the computation of the intersection number, which should be taken zero at the end. The regulators lead to redundancy in both calculation and conclusion.
In this paper, supposing we have constructed $\mathrm{d}\log$ integrand as master integrals, we improve the selection rule of CDE in two aspects. Firstly, we show that CDE matrix is 1-$\mathrm{d}\log$-form coefficients of projecting ($n$+1)-$\mathrm{d}\log$-forms onto $n$-$\mathrm{d}\log$-forms master integrand, while such projecting only need use LO contribution of intersection number. Secondly, we apply the mathematical tool called twisted relative cohomology [43, 44, 45, 46] to avoid regulators, and redundancy as well, at the beginning. Thus we find the exact CDE selection rules.
2 Notations of Intersection Theory and Relative Cohomology
Intersection theory describe integrals take the form $I[u,\varphi]\equiv\int u\,\varphi\,,$ where
$\varphi\equiv\hat{\varphi}(\bm{z})\bigwedge_{j}\mathrm{d}z_{j}=\frac{Q(\bm{z})}{\left(\prod_{k}D_{k}^{a_{k}}\right)\left(\prod_{i}P_{i}^{b_{i}}\right)}\bigwedge_{j}\mathrm{d}z_{j}\,,\quad u=\prod_{i}\left[P_{i}(\bm{z})\right]^{\beta_{i}}\,,\quad a_{k},b_{j}\in\mathbb{N}\,,$ (2)
Feynman integrals and their representation, such as Baikov representation, Feynman parameterization, are integrals in this form.
Define the IBP-equivalence classes of cocycles [41, 42, 47, 48]:
$\bra{\varphi_{L}}\equiv\varphi_{L}\sim\varphi_{L}+\sum_{i}\nabla_{i}\xi_{i}\,,\,\quad\nabla_{i}=dz_{i}\wedge(\partial_{z_{i}}+\hat{\omega}_{i})\,,\quad\hat{\omega}_{i}\equiv\partial_{z_{i}}\log(u).$ (3)
which leads to a cohomology group. Any integral $f_{0}=\int u\varphi_{0}$ (or to be exact, the cocycle $\bra{\varphi_{0}}$ | corresponding to the integrals) in the function family are the vectors in space of cohomology group, and master integrals $f_{i}=\int u\varphi_{i}$ are the basis: $\bra{\varphi_{0}}=\sum_{i}c_{i}\bra{\varphi_{i}}.$ | . $c_{i}$ are the coefficients of reduction and given by projecting, instead of IBP
$c_{I}=\sum_{J}\braket{\varphi_{0}}{\varphi_{J}^{\prime}}\left(\eta^{-1}\right)_{JI}\,,\quad\eta_{IJ}\equiv\braket{\varphi_{I}}{\varphi_{J}^{\prime}}\,,$ (4)
where we use the IBP-invariant inner-product intersection number:
$\braket{\varphi_{L}}{\varphi_{R}}=\sum_{\bm{p}}\operatorname{Res}_{\bm{z}=\bm{p}}\left(\psi_{L}\varphi_{R}\right)\,,\quad\nabla_{1}\cdots\nabla_{n}\psi_{L}=\varphi_{L}\,.$ (5)
In (5), $\bm{p}$ are the points (not lines or hypersurfaces with higher dimension) where n (or more in the case of degenerate pole) hypersurfaces belong to =$\mathcal{B}=\{P_{1}=0,\infty,\cdots,D_{1}=0,\infty,\cdots\}$ meet. ($P_{1}=0,\infty$ means $P_{1}=0,P_{1}=\infty$ ).
Notice that $D_{k}$ in (5) denote the denominators with integer power in $u\varphi$. In the case of Feynman integral, such denominators usually arise from propagators. With them, one could not directly compute (5). One way to handle them is adding regulators of them in u.
$u=\prod_{i}\left[P_{i}(\bm{z})\right]^{\beta_{i}}\prod_{k}D_{k}^{\delta_{k}}$ (6)
Then take the regulators to be zero in the final result. This was adopted in [40] and led to redundancy.
While in this paper, we will adopt the tool called relative cohomology to handle them. We do not go into mathematical details. Roughly speaking, we will let the dual forms $\varphi_{R}$ live on maximal cut of them, and denote them as $\Delta_{\hat{R}}\varphi_{R}$. In practice, it equivalent to applying the maximal cut of the sector of $\varphi_{R}$ to both sides before compute intersection number.
$\braket{\varphi_{L}}{\Delta_{\hat{R}}\varphi_{R}}=\braket{\varphi_{L;\hat{R}}}{\varphi_{R;\hat{R}}}=\sum_{\bm{p}_{\hat{R}}}\operatorname{Res}_{\bm{p}_{\hat{R}}=0}\psi_{L;\hat{R}}\varphi_{R;\hat{R}}\,,\quad\nabla_{1;\hat{R}}\cdots\nabla_{n;\hat{R}}\psi_{L;\hat{R}}=\varphi_{L;\hat{R}}$
$\varphi_{L;\hat{R}}=\operatorname{Res}_{\mathcal{B}_{\hat{R}}}\left(\frac{u\varphi_{L}}{u_{\hat{R}}}\right)\,,\quad\varphi_{R;\hat{R}}=\operatorname{Res}_{\mathcal{B}_{\hat{R}}}\left(\frac{\varphi_{R}u_{\hat{R}}}{u}\right)\,,$
$u_{\hat{R}}=u|_{\mathcal{B}_{\hat{R}}}\,,\quad\mathcal{B}_{\hat{R}}=\{D_{R_{1}}=0,D_{R_{2}}=0,\cdots,D_{R_{n}}=0\}\,.$ (7)
Every symbol with the subscript $\hat{R}$ means it is cut version of the original one. $D_{R_{i}}$ in (7) usually are propagators of the sector of $\varphi_{R}$. If $\mathcal{B}_{\hat{I}}\subseteq\mathcal{B}_{\hat{J}}$, we say $\mathcal{B}_{\hat{I}}$ is a subsector of $\mathcal{B}_{\hat{J}}$.
To keep the information of multivariate pole explicitly in the calculation, we select to analyze and compute intersection number and the multivariate residue in it via multivariate Laurent expansion after factorization transformation corresponding to each poles [48, 40]
$\mathrm{T}^{(\alpha)}:z_{i}\to f^{(\alpha)}_{i}(\bm{x}^{(\alpha)})$
$\mathrm{T}^{(\alpha)}\left[u\right]\equiv u(\mathrm{T}^{(\alpha)}[\bm{z}])=\bar{u}_{\alpha}(\bm{x}^{(\alpha)})\prod_{i}\left[x_{i}^{(\alpha)}-\rho_{i}^{(\alpha)}\right]^{\gamma_{i}^{(\alpha)}}\,.$ (8)
Only after this step, one can apply multivariate expansion legally to $\varphi$ and solve intersection number
$\varphi=\sum_{\bm{b}}\varphi^{(\bm{b})}\,,\quad\varphi^{(\bm{b})}=C^{(\bm{b})}\bigwedge_{i}\left[x_{i}^{(\alpha)}-\rho_{i}^{(\alpha)}\right]^{b_{i}}\mathrm{d}x^{(\alpha)}_{i}\,,$
$\braket{\varphi_{L}}{\varphi_{R}}=\sum_{\alpha}\operatorname{Res}_{\bm{\rho}^{(\alpha)}}\mathrm{T}^{(\alpha)}\left[\psi_{L}\varphi_{R}\right]=\sum_{\alpha}\operatorname{Res}_{\bm{\rho}^{(\alpha)}}\left[\left(\sum_{\bm{b}_{L}}\nabla_{1}^{-1}\cdots\nabla_{n}^{-1}\varphi^{(\bm{b}_{L})}_{L}\right)\sum_{\bm{b}_{R}}\varphi^{(\bm{b}_{R})}_{R}\right]\,.$ (9)
where the powers $\bm{b}=(b_{1},\ldots,b_{n})$.
3 Selection Rule of CDE
3.1 $\mathrm{d}\log$ projection and LO intersection number
For distinction, we use $\mathrm{d}$ to denote the differentiation of integration variables $\bm{z}$ , and use $\hat{\mathrm{d}}$ to denote the differentiation of arbitrary one parameter, such as kinetic parameters or mass. And define $\mathrm{D}=\mathrm{d}+\hat{\mathrm{d}}$.
What we need is analyzing and computing CDE by (4) with (7)
$\big{(}\hat{\mathrm{d}}\Omega\big{)}_{IK}=\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\big{(}\eta^{-1}\big{)}_{JK}\,,\quad\eta_{IJ}\equiv\braket{\varphi_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\,.$ (10)
where $\varphi_{I}$ in original space and dual space are the same $\mathrm{d}\log$ integrand supposed to have been constructed $\varphi_{I}=\bigwedge_{j}\mathrm{d}\log\mathrm{W}^{(I)}_{j}(\bm{z})$. And $\dot{\varphi}_{I}$ in (10) is defined by $\hat{\mathrm{d}}f_{I}\equiv\int u\dot{\varphi}_{I}$. So notice that
$\hat{\mathrm{d}}f_{I}$ $=\int\hat{\mathrm{d}}\left(u\bigwedge_{j}\mathrm{d}\log\mathrm{W}^{(I)}_{j}(\bm{z})\right)=\int u~{}\mathrm{D}\log u\bigwedge_{j}\mathrm{D}\log\mathrm{W}^{(I)}_{j}\,,$ (11)
$\dot{\varphi}_{I}$ is ($n$+1)-$\mathrm{D}\log$-form. Then all we need for computing (10) is formula of LO contribution of intersection number (after factorization of poles)
$\operatorname{Res}_{\bm{\rho}^{(\alpha)}}\mathrm{T}^{(\alpha)}\left[\left(\nabla_{1}^{-1}\cdots\nabla_{n}^{-1}\varphi_{L}\right)\varphi_{R}\right]$ $=\frac{\operatorname{Res}_{\bm{\rho}^{(\alpha)}}\mathrm{T}^{(\alpha)}\left[\varphi_{L}\right]\times\operatorname{Res}_{\bm{\rho}^{(\alpha)}}\mathrm{T}^{(\alpha)}\left[\varphi_{R}\right]}{\prod_{i}\operatorname{Res}_{\rho^{(\alpha)}_{i}}\partial_{x_{i}^{(\alpha)}}\log\left(\mathrm{T}^{(\alpha)}\left[u\right]\right)\mathrm{d}x_{i}^{(\alpha)}}$
$=\frac{C_{L}^{(\bm{b}_{L})}C_{R}^{(\bm{b}_{R})}}{\bm{\gamma}^{(\alpha)}}\,,\quad\quad\bm{\gamma}^{(\alpha)}=\prod_{i}\gamma_{i}^{(\alpha)}$ (12)
3.2 Condition of $\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\neq 0$
Take (12) and that $\varphi_{I}$ is $\mathrm{d}\log$-form into consideration, we could conclude the condition of $\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\neq 0$
$\begin{cases}\mathrm{d}\log\text{-form}&:\ b_{I,i}\geq-1\,,\ b_{J,i}\geq-1\,,\\ \dot{\varphi}_{I;\hat{J}}&:\ b_{\dot{I},j}=b_{I,j}-1\,,\ \text{for one $j$}\,,\\ \braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\neq 0&:\ b_{\dot{I},i}+b_{J,i}\leq-2\end{cases}$
$\Rightarrow\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\neq 0:\ (-2\leq b_{I,j}+b_{J,j}\leq-1)\ \&\ (b_{I,i}=b_{J,i}=-1,i\neq j).$ (13)
We have omit the subscript "$;\hat{J}$" of $b_{*,*}$ for keeping the expression simple, but one should remember they all correspond to signs after maximal cut of sector $\mathcal{B}_{\hat{J}}$. We denote the two case as $\varphi_{I;\hat{J}}$ and $\varphi_{J;\hat{J}}$ share $n_{\hat{J}}$-variable Simple Pole ($n_{\hat{J}}$-SP): $(b_{I,j}+b_{J,j}=-1)\&(b_{I,i}=b_{J,i}=-1,i\neq j)$, and share ($n_{\hat{J}}$-1)-SP: $b_{I,i}=b_{J,i}=-1$.
For the case they share ($n_{\hat{J}}$-1)-SP, $\mathrm{D}\log\left(\mathrm{T}^{(\alpha)}[u]\right)$ in (11) provide pole of one remaining variable, and gives
$\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}=-\frac{\gamma^{(\alpha)}_{j;\hat{J}}}{\bm{\gamma}^{(\alpha)}_{\hat{J}}}\,\hat{\mathrm{d}}\int C_{I}^{(\bm{b}_{I})}C_{J}^{(\bm{b}_{J})}\,\hat{\mathrm{d}}\rho^{(\alpha)}_{j;{\hat{J}}}\,,$ (14)
For the case they share $n_{\hat{J}}$-SP, we have
$\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}=\frac{C_{I}^{(\bm{-1})}C_{J}^{(\bm{-1})}}{\bm{\gamma}^{(\alpha)}_{\hat{J}}}\,\hat{\mathrm{d}}\log\Big{(}\bar{u}_{\alpha;\hat{J}}(\bm{\rho}^{(\alpha)}_{\hat{J}})\Big{)}\,.$ (15)
3.3 Condition of $\big{(}\eta^{-1}\big{)}_{JK}\neq 0$
Notice that $\eta_{JK}$ is non-zero only when $\varphi_{J;\hat{K}}$ and $\varphi_{K;\hat{K}}$ share $n_{\hat{J}}$-SP and each of them gives contribution $C_{J}^{(\bm{-1})}C_{K}^{(\bm{-1})}/\bm{\gamma}^{(\alpha)}_{\hat{K}}$. Also notice that $\big{(}\eta^{-1}\big{)}_{JK}=(-1)^{J+K}|\eta^{(KJ)}|/|\eta|$, in which $\eta^{(KJ)}$ is the adjugate matrix of $\eta$. Then follow the similar analysis as [40], we need modify the conception of "n-SP chain" in [40], which is unoriented, to the following one
  • If $\dot{\varphi}_{I;\hat{J}}$ and $\varphi_{J;\hat{J}}$ share $n_{\hat{J}}$-SP, we say $\varphi_{I}$ is Maximal multivariate Simple Pole (MSP) related to $\varphi_{J}$, and denoted as $\varphi_{I}\rightarrow\varphi_{J}$ or $\varphi_{J}\leftarrow\varphi_{I}$. Now, it is an orient relation. $\varphi_{I}\rightarrow\varphi_{J}$ does not imply $\varphi_{J}\rightarrow\varphi_{I}$. "Maximal" for $\dot{\varphi}_{I;\hat{J}}$ and $\varphi_{J;\hat{J}}$ are $n_{\hat{J}}$-forms and share $n_{\hat{J}}$-SP.
  • If $\varphi_{I}\rightarrow\varphi_{J}$, we also say that $\varphi_{I}$ is linked to $\varphi_{J}$ via a MSP chain. If $\varphi_{I}\rightarrow\varphi_{J}$ and $\varphi_{J}\rightarrow\varphi_{K}$, then we say $\varphi_{I}$ is linked to $\varphi_{K}$ via the MSP chain $\varphi_{I}\rightarrow\varphi_{J}\rightarrow\varphi_{K}$, we denote it as $\varphi_{I}\rightarrow\rightarrow\varphi_{J}$. Similar for more forms.
Then, only when $\varphi_{J}\rightarrow\rightarrow\varphi_{K}$, $\big{(}\eta^{-1}\big{)}_{JK}$ could be non-zero.
3.4 Conclusion
Only when there exist at least a $\varphi_{J}$, satisfy that $\varphi_{J;\hat{J}}$ share at least one $n_{\hat{J}}$-SP or ($n_{\hat{J}}$-1)-SP with $\varphi_{I;\hat{J}}$ and $\varphi_{J}\rightarrow\rightarrow\varphi_{K}$, then $\Omega_{IK}$ could be non-zero. When all $\beta_{i}$ in $u=\prod_{i}\left[P_{i}(\bm{z})\right]^{\beta_{i}}$ are proportional to $\epsilon$, $\Omega_{IK}$ is canonical form.
4 Univariate Example: Compare Regulator with Relative Cohomology
Consider $u=z^{\delta}(z-c_{1})^{\epsilon}(z-c_{2})^{\epsilon}$ with regular $\delta$. Constructing the $\mathrm{d}\log$ basis as
$\varphi_{1}=\frac{\mathrm{d}z}{z}=\mathrm{d}\log z\,,\quad\varphi_{2}=\frac{\mathrm{d}z}{z-c_{1}}-\frac{\mathrm{d}z}{z-c_{2}}=\mathrm{d}\log\left(\frac{z-c_{1}}{z-c_{2}}\right)\,.$ (16)
4.1 Regulator method
If compute the differential equations of the basis via regulator method, we have
$\omega=\mathrm{d}\log u=\left(\frac{\delta}{z}+\frac{\epsilon}{z-c_{1}}+\frac{\epsilon}{z-c_{2}}\right)\mathrm{d}z\,.$
=$\mathcal{B}=\{z=0,z-c_{1}=0,z-c_{2}=0,z=\infty\}\,,\quad\{\bm{p}\}=\{0,c_{1},c_{2},\infty\}\,.$
$\gamma_{1}=\delta\,,\ \ \gamma_{2}=\epsilon\,,\ \ \gamma_{3}=\epsilon\,,\ \ \gamma_{4}=-2\epsilon-\delta\,.$ (17)
Then, $\eta_{11}$ receive contribution from shared 1-SP at $p_{1}=0$ and $p_{4}=\infty$ of $\varphi_{1}$ itself, so it equals $1/\gamma_{1}+1/\gamma_{4}$. Similar analysis also could be applied to other $\eta_{JK}$ and give
$\eta=\braket{\varphi_{I}}{\varphi_{J}}=\left(\begin{array}[]{cc}\frac{1}{\gamma_{1}}+\frac{1}{\gamma_{4}}&0\\ 0&\frac{1}{\gamma_{2}}+\frac{1}{\gamma_{3}}\\ \end{array}\right)=\left(\begin{array}[]{cc}\frac{1}{\delta}-\frac{1}{\delta+2\epsilon}&0\\ 0&\frac{2}{\epsilon}\\ \end{array}\right)\,,$ (22)
$\eta^{-1}=\left(\begin{array}[]{cc}\delta(\delta+2\epsilon)/(2\epsilon)&0\\ 0&\epsilon/2\\ \end{array}\right)\,.$ (25)
For $\braket{\dot{\varphi}_{I}}{\varphi_{J}}$, We have $\dot{\varphi}_{I}$
$\dot{\varphi}_{1}=\mathrm{D}\left(\log\bar{u}_{1}(z)+\delta\log z\right)\wedge\mathrm{D}\log z=\mathrm{D}\log\bar{u}_{1}(z)\wedge\mathrm{D}\log z\,,$
$\dot{\varphi}_{2}=\mathrm{D}\log\bar{u}_{2}(z)\wedge\mathrm{D}\log(z-c_{1})-\mathrm{D}\log\bar{u}_{3}(z)\wedge\mathrm{D}\log(z-c_{2})$
$\quad=\delta~{}\mathrm{D}\log z\wedge\varphi_{2}+2\epsilon~{}\mathrm{D}\log(z-c_{2})\wedge\mathrm{D}\log(z-c_{1})\,,$
$\bar{u}_{1}(z)=(z-c_{1})^{\epsilon}(z-c_{2})^{\epsilon}\,,\quad\bar{u}_{2}(z)=z^{\delta}(z-c_{2})^{\epsilon}\,,\quad\bar{u}_{3}(z)=z^{\delta}(z-c_{1})^{\epsilon}\,.$ (26)
Let’s calculate $\braket{\dot{\varphi}_{1}}{\varphi_{1}}$ as example. For the contribution from the shared 1-SP $p_{1}=0$, one could use (15) and get
$\frac{1}{\delta}\hat{\mathrm{d}}\log\left((z-c_{1})^{\epsilon}(z-c_{2})^{\epsilon}|_{z=0}\right)=\frac{\epsilon}{\delta}\hat{\mathrm{d}}\log(c_{1}c_{2})\,,$ (27)
For the shared 1-SP $p_{4}=\infty$:
$\frac{1}{-\delta-2\epsilon}\hat{\mathrm{d}}\log\left((1-c_{1}t)^{\epsilon}(1-c_{2}t)^{\epsilon}|_{t=0}\right)=\frac{\epsilon}{-\delta-2\epsilon}\hat{\mathrm{d}}\log 1=0\,,$ (28)
where $t=1/z$. Thus it does not contribute. Continue the progress, we will have
$\Big{(}\braket{\dot{\varphi}_{I}}{\varphi_{J}}\Big{)}=\hat{\mathrm{d}}\left(\begin{array}[]{cc}\frac{\epsilon}{\delta}\log\left(c_{1}c_{2}\right)&\log\left(c_{2}/c_{1}\right)\\ \log\left(c_{2}/c_{1}\right)&\frac{\delta}{\epsilon}\log\left(c_{1}c_{2}\right)+4\log(c_{1}-c_{2})\\ \end{array}\right)\,.$ (31)
$\hat{\mathrm{d}}\Omega^{\text{reg}}(c_{i};\epsilon,\delta)=\hat{\mathrm{d}}\left(\begin{array}[]{cc}\frac{\delta+2\epsilon}{2}\log\left(c_{1}c_{2}\right)&\frac{\epsilon}{2}\log\left(\frac{c_{2}}{c_{1}}\right)\\ \frac{\delta(\delta+2\epsilon)}{2\epsilon}\log\left(\frac{c_{2}}{c_{1}}\right)&\frac{\delta}{2}\log\left(c_{1}c_{2}\right)+2\epsilon\log\left(c_{1}-c_{2}\right)\\ \end{array}\right)\,.$ (34)
It is redundant and $\hat{\mathrm{d}}\Omega^{\text{reg}}_{21}$ vanish when $\delta=0$, but the selection rules get from the regulator method in [40] could not tell you that.
4.2 Relative cohomology method
If we calculate the $\Omega$ via relative cohomology method, we have
$u=(z-c_{1})^{\epsilon}(z-c_{2})^{\epsilon}\,,\ \ D_{1}=z$
$\omega=\mathrm{d}\log u=\left(\frac{\epsilon}{z-c_{1}}+\frac{\epsilon}{z-c_{2}}\right)\mathrm{d}z\,.\quad\{\bm{p}\}=\{0,c_{1},c_{2},\infty\}\,.$ (35)
Calculation of $\braket{\varphi_{I}}{\Delta_{\hat{1}}\varphi_{1}}$ will be simplified, become we need apply the cut of $\mathcal{B}_{\hat{1}}$:
$\braket{\varphi_{1}}{\Delta_{\hat{1}}\varphi_{1}}=\braket{\varphi_{1;\hat{1}}}{\varphi_{1;\hat{1}}}=\braket{1}{1}=1$
$\braket{\varphi_{2}}{\Delta_{\hat{1}}\varphi_{1}}=\braket{\varphi_{2;\hat{1}}}{\varphi_{1;\hat{1}}}=\braket{0}{1}=0$
$\braket{\dot{\varphi}_{1}}{\Delta_{\hat{1}}\varphi_{1}}=\braket{\dot{\varphi}_{1;\hat{1}}}{\varphi_{1;\hat{1}}}=\braket{\hat{\mathrm{d}}\log u}{{}_{z=0}}{1}=\epsilon\hat{\mathrm{d}}\log(c_{1}c_{2})$
$~{}\dot{\varphi}_{2}=\mathrm{D}\log u\wedge\mathrm{D}\log\left(z-c_{1}\right)-\mathrm{D}\log u\wedge\mathrm{D}\log\left(z-c_{2}\right)$
$\quad~{}=2\epsilon~{}\mathrm{D}\log\left(z-c_{2}\right)\wedge\mathrm{D}\log\left(z-c_{1}\right)\,,$
$\braket{\dot{\varphi}_{2}}{\Delta_{\hat{1}}\varphi_{1}}=\braket{\dot{\varphi}_{2;\hat{1}}}{\varphi_{1;\hat{1}}}=\braket{0}{1}=0\,.$ (36)
It follows that
$\eta=\braket{\varphi_{I;\hat{J}}}{\varphi_{J;\hat{J}}}=\left(\begin{array}[]{cc}1&0\\ 0&2/\epsilon\\ \end{array}\right)\,,\quad\eta^{-1}=\left(\begin{array}[]{cc}1&0\\ 0&\epsilon/2\\ \end{array}\right)\,,$ (41)
$\Big{(}\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\Big{)}=\hat{\mathrm{d}}\left(\begin{array}[]{cc}\epsilon\log\left(c_{1}c_{2}\right)&\epsilon\log\left(c_{2}/c_{1}\right)\\ 0&4\log\left(c_{1}-c_{2}\right)\\ \end{array}\right)\,$ (44)
$\hat{\mathrm{d}}\Omega=\Big{(}\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\Big{)}.\eta^{-1}=\epsilon~{}\hat{\mathrm{d}}\left(\begin{array}[]{cc}\log\left(c_{1}c_{2}\right)&\frac{1}{2}\log\left(c_{2}/c_{1}\right)\\ 0&2\log\left(c_{1}-c_{2}\right)\\ \end{array}\right)\,.$ (47)
We have $\hat{\mathrm{d}}\Omega=\hat{\mathrm{d}}\Omega^{\text{reg}}(c_{i};\epsilon,0)$. It shows that the relative cohomology method simplifies the computation, avoids redundancy of regulators, and gets the exact selection rules of CDE as it could tell you $\Omega_{21}=0$ here.
5 Summary and Outlook
We improve the selection rules of CDE, provide the $\mathrm{d}\log$ projection perspective, and avoid redundancy via relative cohomology. The mathematical structure we have found may benefit many aspects in evaluating perturbation quantum field theory, such as reduction (since CDE themselves are reduction relations), symbology bootstrap, and analytic solution of Feynman integrals.
Acknowledgements
We thank Andrzej Pokraka for valuable discussions. This work is supported by Chinese NSF funding under Grant No.11935013, No.11947301, No.12047502 (Peng Huanwu Center), No.12247120, No.12247103, NSAF grant No.U2230402, and China Postdoctoral Science Foundation No.2022M720386.